OpenStudy (anonymous):

The vector has a magnitude of 10. The x- and y-components of is positive and equal and its z-component is 4. Determine x and y component of vector .

OpenStudy (anonymous):

\[|\vec{u}| = 10\] \[\implies \sqrt{u_x^2 + u_y^2 + u_z^2}=10\]\[\implies u_x^2 + u_y^2 + u_z^2 = 100\] \[u_z = 4\]\[u_x = u_y\] Solve for \(u_x\) and \(u_y\) (since they're the same)